Im asking for help and the solution to how to find this answer. I would like to now how you got it, not just the answer. Thank you to all who are whiling..
If event A is a proper subset of event B, what is p(A|B)?

Answer by rothauserc(4718)

the definition of P(A|B) is P(A and B) / P(B),
now given event A is a proper subset of event B, then
*****************************************************
P(A|B) = P(A) / P(B) for A contained in B
P(A|B) = P(0) / P(B) = 0 for A = empty set